# Periodic Vibration Gaussian: Dynamic Urban Scene Reconstruction and Real-time Rendering

### v2

@@ 4.1 Comparison with state of the art
Results on KITTI Our approach is also quantitatively evaluated on the KITTI benchmark, following SUDS [35]. We select sequences characterized by extensive movement for analysis. The proposed method surpasses all competitors across every evaluated metric. Notably, while these sequences present a substantial number of dynamic objects, our temporal smoothing mechanism secures a concise scene representation, thereby mitigating over-fitting and ensuring superior image quality in novel viewpoints. Table 1 and Fig. 5 demonstrate our significant improvement over the leading SUDS [35] in image reconstruction, with improvements of $16.0\%$ in PSNR, $7.0\%$ in SSIM, and a $62.2\%$ decrease in LPIPS and surpass the leading EmerNeRF [45] in novel view synthesis with improvement of $8.7\%$ in PSNR, $11.9\%$ in SSIM, and a $51.9\%$ decrease in LPIPS. 3DGS achieves much inferior results on KITTI as there were much more dynamic objects in the scenes than Waymo.
@@ 4.2 Ablation study
Table 5: Ablation on novel view synthesis.
Figure 9: Temporal smoothing (a) on and (b) off.
Refer to caption (a) On Refer to caption (b) Off
Refer to caption (a) Constant Refer to caption (b) Linear Refer to caption (c) Ours
Figure 9: Temporal smoothing (a) on and (b) off.
(a) Effect of the depth $\mathcal{L}_{d}$ and sky module.
(c) Effect of Positional-Aware Control (PAC) in improving the depth of distant components.
Figure 11: More visualization of the ablation study.
In Table 5, we conduct an ablation study to investigate the impact of the primary components of our model on novel view synthesis within dynamic scenes, specifically utilizing the Waymo Open Dataset. We set $\eta=1$ to deactivate the scene flow-based temporal smoothing mechanism, a crucial element in RGB rendering. As illustrated in Fig. 10, the temporal smoothing mechanism results in enhanced smoothness in novel view rendering and promotes temporal and spatial consistency of PVG points. Meanwhile, temporal smoothing can help the model to learn the correct motion trend as shown in Fig. 10(b). Our findings indicate that the integration of LiDAR supervision, sky refinement module contributes to more plausible geometry (see in Fig. 10(a)). Depth loss and sky module make a minimal effect on novel view RGB rendering mainly because test novel views are interpolations of training views (both in the driving paths), making these metrics not fully reflect the quality of synthesis. Fig. 10(c) demonstrates positional-aware control strategy significantly improves the reconstruction of distant views. Additionally, Fig. 10(b) shows the inclusion of velocity loss facilitates the convergence of velocity to a sparser rank, thereby simplifying the segmentation process between dynamic and static elements.
@@
Dynamics formulation We investigate the influence of $\widetilde{\bm{\mu}}(t)$ formulation. We compare our design (Eq. (6)) with two alternatives: (a) A constant model, i.e., $\widetilde{\bm{\mu}}(t)=\bm{\mu}$, and (b) A linear model, i.e., $\widetilde{\bm{\mu}}(t)=\bm{\mu}+\bm{v}(t-\tau)$. We observe from Fig. 10 that (1) The constant model adeptly represents static elements (see yellow box) but exhibits suboptimal performance on dynamic elements (red box); (2) The linear model can more smoothly capture dynamic aspects. However, it consistently misrepresents static components resulting in ambiguities and ghosting effects. Our method overcomes both limitations by using a sine function based vibration design.
5 Conclusion
@@
Appendix 0.A More model interpretation
@@
Due to sparse training frames, the renderings of novel timestamps are prone to underfitting. To make the $\{\mathcal{H}_{i}\}$ temporally and spatially consistent with its intrinsic motion, we exploit the inherent temporal consistency law: The status $\mathcal{H}_{i}(t)$ at time $t$ is similar to the result ^$\widehat{\mathcal{H}}_{i}(t)$ of translating the status to $t-\Delta t$ by a distance $\bm{v}\Delta t$. This introduces an additional optimization regularization defined as:
@@
Appendix 0.B More implementation details
@@
Appendix 0.C Velocity map and staticness map
(a) render RGB
(b) Velocity map
(c) Staticness map
We visualize the renderings of velocity map $\mathcal{\bar{V}}$ and staticness $\rho$ map to analyze the behavior of PVG. For the visualization of $\mathcal{\bar{V}}$, we first transform $\bm{\bar{v}}$ of each pixel from the world coordinate system to the camera coordinate system and project the $\bm{v}_{cam}$ onto a plane parallel to the camera plane. Then we use the color coding of optical flow for visualization.
For the visualization of staticness $\rho$ map, we need to clamp each point’s $\rho$ to the range of $[0,2]$ and render, otherwise the visualization is not visually distinctive.
As shown in Fig. 14, our model captures not only dynamic objects like cars and people, but also moving light and shadow. Note that, pixels with small $\rho$ point to dynamic areas and large $\rho$ for static ones.
Appendix 0.D Visualization for Waymo
@@
We also test the rendering quality of PVG under different camera settings, for evaluating the robustness of novel view synthesis. As in Fig. 15, we zoom in/out, disturb the camera extrinsics and intrinsics for novel view synthesis, and PVG remain stable and compact rendering quality. $\mathcal{L}_{d}$ improves geometry and (e.g., road and dynamics) and increases generalization performance beyond the driving path (Fig. 14(d) vs. 14(e)).
(a) Origin
(b) Zoom in
(c) Zoom out
(d) Shift right
(e) w/o $\mathcal{L}_{d}$
Figure 15: Rendered images under different camera settings.
Appendix 0.E Visualization for KITTI
@@
Appendix 0.F Results on small-scale data
Table 6: Per-scene results on the Plenoptic Video dataset.
Coffee Martini Cook Spinach Cut Beef Flame Salmon Flame Steak Sear Steak Mean
HexPlane [7] - 32.04 32.55 29.47 32.08 32.39 31.70
K-Planes-explicit [12] 28.74 32.19 31.93 28.71 31.80 31.89 30.88
K-Planes-hybrid [12] 29.99 32.60 31.82 30.44 32.38 32.52 31.63
MixVoxels [37] 29.36 31.61 31.30 29.92 31.21 31.43 30.80
NeRFPlayer [29] 31.53 30.56 29.353 31.65 31.93 29.12 30.69
HyperReel [2] 28.37 32.30 32.922 28.26 32.20 32.57 31.10
PVG (ours) 28.64 32.85 33.14 29.65 33.45 33.43 31.86
For further evaluate the capability of our model, we evaluate PVG on a small-scale benchmark Plenoptic Video for dynamic 3D reconstruction. As shown in Table 6, our model outperforms previous methods which are particularly designed for small-scale dynamic reconstruction by a clear margin.
Appendix 0.G Limitation
Unlike neural network based representation models, our PVG uses independent and discrete points to represent a scene. Despite the advantages including high flexibility, simple composition, strong expression ability, and strong fitting ability, the independence of each point makes it more difficult to model the smoothness over time and space. Our temporal smoothing mechanism does enhance the correlation between points, but still not fully solves this problem. More fine-grained, dedicated designs are needed for further enhancement.
@@ Appendix 0.G Limitation
S-NeRF [43] Refer to caption
StreetSurf [14] Refer to caption
3DGS [15] Refer to caption
NSG [23] Refer to caption
SUDS [35] Refer to caption
Mars [41] Refer to caption
EmerNeRF [45] Refer to caption
PVG (Ours) Refer to caption
GT Refer to caption
Figure 18: Qualitative results of image reconstruction on KITTI.
S-NeRF [43] Refer to caption
StreetSurf [14] Refer to caption
3DGS [15] Refer to caption
NSG [23] Refer to caption
SUDS [35] Refer to caption
Mars [41] Refer to caption
EmerNeRF [45] Refer to caption
PVG (Ours) Refer to caption
GT Refer to caption
Figure 19: Qualitative results of novel view synthesis on KITTI.

### v3

@@ 4.1 Comparison with state of the art
Results on KITTI Our approach is also quantitatively evaluated on the KITTI benchmark, following SUDS [9]. We select sequences characterized by extensive movement for analysis. The proposed method surpasses all competitors across every evaluated metric. Notably, while these sequences present a substantial number of dynamic objects, our temporal smoothing mechanism secures a concise scene representation, thereby mitigating over-fitting and ensuring superior image quality in novel viewpoints. Table 1 and Fig. 4 demonstrate our significant improvement over the leading SUDS [9] in image reconstruction, with improvements of $16.0\%$ in PSNR, $7.0\%$ in SSIM, and a $62.2\%$ decrease in LPIPS and surpass the leading EmerNeRF [10] in novel view synthesis with improvement of $8.7\%$ in PSNR, $11.9\%$ in SSIM, and a $51.9\%$ decrease in LPIPS. More visualization results can be seen in Fig. 17 and 18. 3DGS achieves much inferior results on KITTI as there were much more dynamic objects in the scenes than Waymo.
(a) DeformableGS [49] (b) 4DGS [37] (c) StreetGS [44] (d) OmniRe [45] (e) PVG (Ours) (f) GT
Figure 8: Novel view synthesis of two scenes on the Waymo dataset.
Table 4: Comparison with Gaussian Splatting models on the Waymo dataset. We report average PSNR, SSIM, and LPIPS for both full images and dynamic vehicle regions. For the pixel-level metric PSNR, results are also separated for static (S) and dynamic (D) parts of the scene. Following OmniRe [45], all evaluations use a resolution of 960×640. The running speed is tested using a NVIDIA RTX A6000 GPU. Box: The need for object bounding boxes for model training.
Image reconstruction Novel view synthesis
[…]
StreetGaussian [44] Jan 24 67.8 32.18 0.918 0.090 32.68 28.64 28.92 0.877 0.110 29.38 25.54
HUGS [47] Mar 24 46.1 31.58 0.904 0.094 32.79 25.78 29.34 0.865 0.110 30.42 23.84
Omnire [45] Aug 24 58.3 32.32 0.923 0.084 32.92 28.36 29.41 0.884 0.101 29.91 25.85
4DGS [37] $\times$ Oct 23 67.9 34.94 0.935 0.068 35.36 31.80 26.56 0.793 0.132 26.87 23.97
DeformableGS [49] $\times$ Sep 23 4.9 32.34 0.923 0.086 32.94 27.39 29.52 0.889 0.100 30.37 24.66
PVG (Ours) $\times$ Nov 23 70.4 36.02 0.952 0.059 36.79 32.17 30.95 0.897 0.105 31.72 25.82
4.2 Comparison with concurrent Gaussian Splatting models
We next conduct an extensive comparison against concurrent 3D Gaussian Splatting methods. For this experiment, we evaluated 8 scenes from StreetGaussian [44] and 4 original scenes from our paper, with each scene captured using 3 cameras at a resolution of 960×640. The results are reported in Table 4 and Figure 8.
Existing methods can be broadly categorized into two groups. The first are box-based methods (e.g., StreetGaussian, HUGS, Omnire), which rely on object-level supervision from 3D bounding boxes. While this approach can achieve high PSNR scores for dynamic objects, it makes the model more complex and less flexible. This reliance on automatically generated bounding boxes can introduce noisy supervision and can struggle with distant or occluded objects, where detection performance is poor. Furthermore, treating dynamic objects separately from the background can reduce scene consistency.
The second group are box-free methods (e.g., 4DGS, DeformableGS), which are more principled as they learn scene dynamics directly without external supervision. However, intuitive strategies like extending 3DGS to spatiotemporal 4D Gaussian primitives (4DGS) or learning a deformation field (DeformableGS) are often inferior when dealing with the fast motion and sparse views typical of urban scenes. As our evaluation confirms, PVG achieves the best overall performance among all competitors, without reliance on object bounding boxes. The results show our superior ability to model dynamic scenes, particularly with challenging distant and occluded objects. This is because, our PVG model quantifies dynamics at the primitive level with meaningful concepts like life peak and cycle, allowing it to model both static and dynamic elements in a unified manner. This approach avoids the limitations of both box-based methods and other box-free methods.
(a) render RGB
(b) Velocity map
(c) Staticness map
Full    Static    Dynamic
Figure 10: Scene separation into static and dynamic elements by PVG.
4.3 Dynamic element analysis
We visualize the renderings of velocity map $\mathcal{\bar{V}}$ and staticness $\rho$ map to analyze the behavior of PVG. For the visualization of $\mathcal{\bar{V}}$, we first transform $\bm{\bar{v}}$ of each pixel from the world coordinate system to the camera coordinate system and project the $\bm{v}_{cam}$ onto a plane parallel to the camera plane. Then we use the color coding of optical flow for visualization. For the visualization of staticness $\rho$ map, we need to clamp each point’s $\rho$ to the range of $[0,2]$ and render, otherwise the visualization is not visually distinctive. Note, pixels with small $\rho$ point to dynamic areas and large $\rho$ for static ones.
As shown in Fig. 9, PVG captures not only dynamic objects like cars and people, but also moving light and shadow. We see from Figure 10 that, static regions remain visually stable and robust to temporal variations, while dynamic regions such as vehicles and trees are accurately captured. This demonstrates PVG’s inherent ability to effectively separate and render static and dynamic scene elements.
@@ 4.4 Ablation study
We conduct an ablation study to investigate the impact of the primary components of PVG on novel view synthesis of dynamic scenes on the Waymo Open Dataset. We set $\eta=1$ to deactivate the scene flow-based temporal smoothing mechanism, a crucial element in RGB rendering. The results are given in Table 6. The temporal smoothing mechanism results in enhanced smoothness in novel view rendering and promotes temporal and spatial consistency of PVG points (Fig. 12). Meanwhile, temporal smoothing can help the model to learn the correct motion trend (Fig. 13(b)). Our findings indicate that the integration of LiDAR supervision, sky refinement module contributes to more plausible geometry (Fig. 13(a)). Depth loss and sky module make a minimal effect on novel view RGB rendering mainly because test novel views are interpolations of training views (both in the driving paths), making these metrics not fully reflect the quality of synthesis. Fig. 13(c) demonstrates positional-aware control strategy significantly improves the reconstruction of distant views. Additionally, Fig. 13(b) shows the inclusion of velocity loss facilitates the convergence of velocity to a sparser rank, thereby simplifying the segmentation process between dynamic and static elements.
(a) On
(b) Off
Figure 11: Temporal smoothing (a) on and (b) off.
(c) Constant
(d) Linear
(e) Ours
(a) Effect of the depth $\mathcal{L}_{d}$ and sky module.
(c) Effect of Positional-Aware Control (PAC) in helping the depth of distant components.
Figure 13: More visualization of the ablation study.
@@
Dynamics formulation We investigate the influence of $\widetilde{\bm{\mu}}(t)$ formulation. We compare our design (Eq. (6)) with two alternatives: (a) A constant model, i.e., $\widetilde{\bm{\mu}}(t)=\bm{\mu}$, and (b) A linear model, i.e., $\widetilde{\bm{\mu}}(t)=\bm{\mu}+\bm{v}(t-\tau)$. We observe from Fig. 12 that (1) The constant model adeptly represents static elements (see yellow box) but exhibits suboptimal performance on dynamic elements (red box); (2) The linear model can more smoothly capture dynamic aspects. However, it consistently misrepresents static components resulting in ambiguities and ghosting effects. This is because linear model has a hard time optimizing static points to the right position and zero speed. Our method overcomes both limitations by using a sine function based vibration design.
[…]
$l$ PSNR↑ SSIM↑ LPIPS↓
constant 27.77 0.848 0.281
0.1 27.83 0.845 0.278
0.2 28.11 0.849 0.279
0.5 28.10 0.846 0.283
1.0 28.01 0.843 0.289
linear 27.09 0.841 0.296
Table 8: Ablation on the scale factor threshold $r$. Off means we do not use the positional-aware control strategy.
$r$ PSNR↑ SSIM↑ LPIPS↓
[…]
$\infty$/Off 27.90 0.846 0.282
Table 9: Ablation on temporal smoothing probability $\eta$.
Image reconstruction Novel view synthesis
[…]
0.25 33.08 0.914 0.215 27.67 0.837 0.288
0.50 32.46 0.910 0.229 28.11 0.849 0.279
0.75 32.60 0.909 0.224 28.16 0.851 0.276
1.00 32.27 0.905 0.230 28.18 0.852 0.276
Effect of cycle length $l$ We investigated how different values of $l$ affect the reconstruction quality. The setting $l=0$ corresponds to the constant case, while $l\to\infty$ corresponds to the linear case. As shown in Table 8, the best overall performance is achieved when $l=0.2$. We observed that smaller values of $l$ lead to finer reconstruction of static elements, but overly small values can negatively affect the reconstruction of dynamic elements. A moderately chosen $l$ strikes a good balance between static and dynamic reconstruction quality, leading to the best overall results—as illustrated in Fig. 12. Meanwhile, the reconstruction quality remains stable across a wide range of $l$ values, as long as extreme value are avoided.
Effect of scale factor $r$ We tested the effect of varying the threshold $r$ on reconstruction quality. Specifically, we define the base radius $r_{0}$ as the ego vehicle’s travel range and scale it by different factors to assess its influence. As shown in Table 8, as long as this strategy is applied, the reconstruction performance remains largely consistent across different values of $r$. This is because the threshold mainly affects distant background regions, which are generally less sensitive to the choice of $r$. As a result, PVG demonstrates strong robustness with respect to this parameter.
Effect of temporal smoothing probability $\eta$ We evaluated the influence of $\eta$ on both reconstruction and novel view synthesis performance. As reported in Table 9, setting $\eta$ to a smaller value notably degrades the quality of novel view synthesis. This is primarily because weaker temporal regularization leads to insufficient self-supervision, impairing the model’s generalization capability. Conversely, increasing $\eta$ introduces stronger temporal constraints, which may slightly compromise reconstruction accuracy due to potential over-regularization. However, beyond a certain threshold, the performance on both tasks plateaus, indicating that the model is relatively robust to the choice of $\eta$ within a broad range. These results highlight the importance of temporal smoothing while demonstrating the stability of PVG.
4.5 View synthesis of different camera settings
We test the rendering quality of PVG under different camera settings, for evaluating the robustness of novel view synthesis. As in Fig. 14, we zoom in/out, disturb the camera extrinsics and intrinsics for novel view synthesis, and PVG remains stable and compact rendering quality. $\mathcal{L}_{d}$ improves geometry and (e.g., road and dynamics) and increases generalization performance beyond the driving path (Fig. 14(d) vs. 14(e)).
(a) Origin
(b) Zoom in
(c) Zoom out
(d) Shift right
(e) w/o $\mathcal{L}_{d}$
Figure 14: Rendered images under different camera settings.
4.6 Evaluation on Plenoptic video dataset
Table 10: Per-scene PSNR results on the Plenoptic video dataset.
Coffee Martini Cook Spinach Cut Beef Flame Salmon Flame Steak Sear Steak
4DGS(K-Planes) [35] 27.34 32.46 32.90 29.20 32.51 32.49
4DGS [51] 28.33 32.93 33.85 29.38 34.03 33.51
PVG (Ours) 28.64 33.01 33.20 29.65 33.45 33.60
To further assess the effectiveness of our model, we conduct experiment on the Plenoptic video dataset [24] which was widely evaluated in those 4D Gaussian splatting methods [35, 37]. This dataset consists of six real-world scenes of approximately ten seconds each and involves diverse human motions. For each scene, one view is held out for testing, while the remaining views are used for training. To ensure consistency, the reported baselines [35, 37] are directly taken from their original papers. As shown in Table 10 above, PVG achieves competitive performance, particularly in dynamic scenes with complex motion and background such as water flow.
5 Conclusions
@@
Limitations Our PVG is built upon independent and discrete Gaussian points. While this design brings advantages such as flexibility, compositionality, and strong fitting capacity, it also introduces challenges in enforcing spatial and temporal coherence for scenarios involving fast-moving dynamic objects. Additionally, due to the independent nature of the points, efficiently editing dynamic objects remains difficult. We consider addressing these challenges through more fine-grained motion modeling and structured priors in future work.
6 Data availability statement
The datasets generated and/or analysed during the current study are available in the Waymo Open Dataset [1] (https://waymo.com/open), KITTI [2] (https://www.cvlibs.net/datasets/kitti/index.php) and Plenoptic video dataset [24] (https://github.com/facebookresearch/Neural_3D_Video/releases/tag/v1.0).
Appendix A More model interpretation
@@
Due to sparse training frames, the renderings of novel timestamps are prone to underfitting. To make the $\{\mathcal{H}_{i}\}$ temporally and spatially consistent with its intrinsic motion, we exploit the inherent temporal consistency law: The status $\mathcal{H}_{i}(t)$ at time $t$ is similar to the result $\widehat{\mathcal{H}}_{i}(t)$ of translating the status to $t-\Delta t$ by a distance $\bm{v}\Delta t$. This introduces an additional optimization regularization defined as:
@@
Appendix B More implementation details
@@
Appendix C Visualization for Waymo
@@
Appendix D Visualization for KITTI
@@
S-NeRF [5] Refer to caption
StreetSurf [6] Refer to caption
3DGS [11] Refer to caption
NSG [7] Refer to caption
SUDS [9] Refer to caption
Mars [43] Refer to caption
EmerNeRF [10] Refer to caption
PVG (Ours) Refer to caption
GT Refer to caption
Figure 17: Qualitative results of image reconstruction on KITTI.
S-NeRF [5] Refer to caption
StreetSurf [6] Refer to caption
3DGS [11] Refer to caption
NSG [7] Refer to caption
SUDS [9] Refer to caption
Mars [43] Refer to caption
EmerNeRF [10] Refer to caption
PVG (Ours) Refer to caption
GT Refer to caption
Figure 18: Qualitative results of novel view synthesis on KITTI.